OpenStudy (anonymous):

Use the properties of addition and multiplication to find the missing number. (And what property is this) 5x = x·___

jimthompson5910 (jim_thompson5910):

You can multiply any two numbers in any order for example 2*3 = 6 3*2 = 6 or 7*9 = 63 9*7 = 63

jimthompson5910 (jim_thompson5910):

this rule also applies to variables because variables are just placeholders for numbers that means 5*x = x*5 this rule is formally called the commutative property of multiplication http://www.aaamath.com/pro74bx2.htm